OpenStudy (anonymous):

Find the equation of the line with the given properties. Express the equation in general form or slope-intercept form. perpendicular to the line 2x+y=3; contains the point (-2,6) The equation of the line is____________ Simplify your answer. Type answer in general form or slope-intercept form.

OpenStudy (anonymous):

start by solving for y

OpenStudy (calculusfunctions):

If given the equation of a line in the form Ax + By = C, then the slope of the line perpendicular to this line is given by\[m =\frac{ B }{ A }\]

OpenStudy (anonymous):

y=2x+3 y-6=1/2x+2 I got this part

OpenStudy (anonymous):

slow down for a minute

OpenStudy (calculusfunctions):

Try again.

OpenStudy (anonymous):

first you need to find the inverse of your equation The gradient of the new line is -(1/grad of old line).

OpenStudy (anonymous):

gradient=slope

OpenStudy (anonymous):

@malibugranprix2000 do you understand or do you need me to explain it more?

OpenStudy (anonymous):

I think I need help, please.

OpenStudy (anonymous):

ok so how would you get y by itself?

OpenStudy (anonymous):

by moving the 2x+3

OpenStudy (anonymous):

but if 2x is positive you need to subtract it to the other side so you would get y=-2x+3

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so what is your slope in this equation?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ok now the slope of the new line is -(1/slope)

OpenStudy (anonymous):

just let me know if you dont understand where i am going and i will help you through it

OpenStudy (anonymous):

so what is the slope?

OpenStudy (anonymous):

the slope of the new line is -(1/(-2))

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

so that simplified turns out to be what?

OpenStudy (anonymous):

is it 1/2

OpenStudy (anonymous):

ok so if we put that into the equation we get y=(1/2)x+b correct

OpenStudy (anonymous):

right

OpenStudy (anonymous):

in the question they also gave you an x and y value plug those into the equation

OpenStudy (anonymous):

do you understand how to do that?

OpenStudy (anonymous):

@malibugranprix2000 ?

OpenStudy (anonymous):

that is plugged in as x1 and y1

OpenStudy (anonymous):

plug it into the equation i just gave you

OpenStudy (anonymous):

6=-2x+3

OpenStudy (anonymous):

leave b alone and change x instead

OpenStudy (anonymous):

and we are entering that into y=(1/2)x+b this equation

OpenStudy (anonymous):

@malibugranprix2000 come on i want to get to bed not trying to be mean or anything

OpenStudy (anonymous):

oh I am sorry you can go thats fine.

OpenStudy (anonymous):

no i want to help we can get it done in 5 min i you just bear with me

OpenStudy (anonymous):

so we have 6=(1/2)(-2)+b right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

6=-1+b

OpenStudy (anonymous):

ok now get b by itself

OpenStudy (anonymous):

so I am not sure b=-1+6?

OpenStudy (anonymous):

you need to add one to both sides

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

6+1=-1+6+1

OpenStudy (anonymous):

because you are adding the same thing to both side they cancel out meaning that the equation is still equal

OpenStudy (anonymous):

oh so its 7=-5+1

OpenStudy (anonymous):

sorry 6+1=-1+b+1

OpenStudy (anonymous):

so the -1 and 1 cancel out and leave b by itself

OpenStudy (anonymous):

7=-1+b

OpenStudy (anonymous):

nope its just 7=b

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

understand how that works?

OpenStudy (anonymous):

yes, thanks a lot

OpenStudy (anonymous):

now take the b that we just found and plug it into our new equation y=(1/2)x+b

OpenStudy (anonymous):

y=(1/2)x+7

OpenStudy (anonymous):

thats it :) hope i helped

OpenStudy (anonymous):

let me know if you dont understand any part of it

OpenStudy (anonymous):

I think I got it, thanks for your help.

OpenStudy (anonymous):

np